# On spectrum of corona product of duplication signed graph and its application

### v4

@@ 4.2 Equienergetic signed graphs
In the following Theorem 4.3 we describe a method for constructing non-co-spectral equienergetic signed graphs.
@@
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two non-co-spectral equienergetic signed graphs of order $m$ with same coronal, then for any arbitrary signed graph $\Gamma$ of order $n$, the signed graphs $\Gamma\circledast\Gamma_{1}$ and $\Gamma\circledast\Gamma_{2}$ are also non-co-spectral equienergetic.
@@
The above theorem can be used to generate families of non-co-spectral equienergetic signed graphs from a given pair of non-co-spectral equienergetic signed graphs with same coronals. The following corollary gives the existence of such non trivial non-co-spectral equienergetic signed graphs with same coronals.
@@
Suppose that $\Gamma_{1}$ and $\Gamma_{2}$ are non-co-spectral equienergetic $r$-regular signed graphs. It is known that the coronal of any two $r$-regular signed graphs are equal. Hence by Theorem 4.3, the signed graphs $\Gamma\circledast\Gamma_{1}$ and $\Gamma\circledast\Gamma_{2}$ are non-co-spectral energetic signed graphs. ∎
Acknowledgement
We would like to acknowledge National Institute of Technology Sikkim for giving doctoral fellowship to Bishal Sonar.
@@ Acknowledgement
Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.

### v5

@@ 4.2 Equienergetic signed graphs
In the following Theorem 4.3, we describe a method for constructing non-co-spectral equienergetic signed graphs.
@@
Let $\Gamma_{1}$ and $\Gamma_{2}$ be two non-co-spectral equienergetic signed graphs of order $m$ with the same coronal, then for any arbitrary signed graph $\Gamma$ of order $n$, the signed graphs $\Gamma\circledast\Gamma_{1}$ and $\Gamma\circledast\Gamma_{2}$ are also non-co-spectral equienergetic.
@@
The above theorem can generate families of non-co-spectral equienergetic signed graphs from a given pair of non-co-spectral equienergetic signed graphs with the same coronals. The following corollary gives the existence of such non-trivial non-co-spectral equienergetic signed graphs with the same coronals.
@@
Suppose that $\Gamma_{1}$ and $\Gamma_{2}$ are non-co-spectral equienergetic $r$-regular signed graphs. The coronal of any two $r$-regular signed graphs is known to be equal. Hence by Theorem 4.3, the signed graphs $\Gamma\circledast\Gamma_{1}$ and $\Gamma\circledast\Gamma_{2}$ are non-co-spectral energetic signed graphs. ∎
5 Conclusion
These results establish a spectral framework for analyzing the interplay between the structural properties of the base signed graphs and their corona products. As an application, we showcase the practical utility of the duplication-signed corona product in generating integral and equienergetic signed graphs. These graphs are of great interest in graph theory because of their combinatorial and spectral significance. We give concrete examples to demonstrate the construction and verification of these properties. This work expands the study of signed graphs and their spectral properties, providing fresh opportunities to explore concepts like structural balance and graph spectra. It also shows how these ideas can be applied to create specialized graph classes, making meaningful contributions to the field of graph theory.
Acknowledgment
We want to acknowledge the National Institute of Technology Sikkim for giving a doctoral fellowship to Bishal Sonar and Satyam Guragain.
@@ Acknowledgment
Data sharing does not apply to this article as no datasets were generated or analyzed during the current study.